### v2

@@ 3.1 General beliefs
In this subsection, we suppose $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ are general probability measures. We first introduce the following notation. Let $\Pi_{n}(\Omega)=\{(A_{1},\dots,A_{n}):\cup_{i=1}^{n}A_{i}=\Omega,~{}A_{i}\cap A_{j}=\emptyset,~{}i\neq j\}$. For $\Lambda_{1},\dots,\Lambda_{n}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, let
$\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=\inf\left\{\sum_{i=1}^{n}y_{i}:\mathbb{Q}_{i}\left(X>\sum_{i=1}^{n}y_{i},A_{i}\right)\leqslant\Lambda_{i}(y_{i})~{}\text{for some}~{}(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$
@@
For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, we have
@@
$X_{i}=\left(X-\sum_{i=1}^{n}y_{i}^{*}\right)\mathds{1}_{A_{i}^{*}}+y_{i}^{*},~{}i=1,\dots,n.$ (3)
@@
Note that Theorem 1 is very general, covering Theorems 1, 2, 4 of Liu (2024) and Theorem 4 of Embrechts et al. (2020) as special cases. Compared with the results in Liu (2024), our results in Theorem 1 offer semi-explicit formulas for the inf-convolution. Later, we will specify the relation of $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ to obtain more explicit formulas for the inf-convolution. Moreover, Theorem 1 does not explicitly show when the inf-convolution is $\infty$- ∞. The discussion on the finiteness of $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$ is complicated as it heavily relies on the properties of Lambda functions and the relationship between $\mathbb{Q}_{i}$; see Liu (2024) and Proposition 1 in Section 2. This will be discussed in the next two propositions. Finally, note that if all $\Lambda_{i}$ are constants, then Theorem 1 boils down to Theorem 4 of Embrechts et al. (2020), where the inf-convolution of $\mathrm{VaR}$ has been studied with belief heterogeneity.
@@
Let $\mathcal{X}=L^{\infty}$. For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, we have
@@
    If $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$ for all $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>-\infty$.
Note that if all $\Lambda_{i}$ are constants, then Proposition 2 offers a sufficient and necessary condition for the finiteness of the inf-convolution. In general, due to the complexity of the $\Lambda_{i}$ functions, those conditions are sufficient but not necessary.
@@
Let $\mathcal{X}=L^{\infty}$. Suppose all $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1 are decreasing. Then we have the following conclusion.
@@
    If $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}<1$ for some $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=-\infty$;
@@
    If $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$ for all $(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>-\infty$.
Suppose all $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1 are increasing. Then we have the following conclusion.
@@
In this more refined result, we see that $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}$ plays a pivotal role when $\Lambda$ functions are decreasing. We see that, in order to guarantee a fininte inf-convolution for decreasing $\Lambda_{i}$, it is necessary for this quantity to exceed 1. For simplicity, assume that for all $i$, $\lambda_{i}^{+}$ is attained by $\Lambda_{i}$. Then, the condition in case (ii) of Proposition 3 for deceasing $\Lambda_{i}$ is violated when one can find a partition of $\Omega$ such that for all agents we have $\mathbb{Q}_{i}(A_{i})\leqslant\lambda_{i}^{+}$. This means that for each agent we can find a set $A_{i}$ on which the risk is suitably bounded by the largest acceptable probability to the agent of an adverse tail event. Motivated by this, we may consider $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}$ as a measure of belief homogeneity, which also reflects preferences. Consider two extreme cases. First, if sets $A_{1},\dots,A_{n}$ exist for which we have $\mathbb{Q}_{i}(A_{i})=0$ for all $i$, then we will have $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}=0.$ The ability to find such sets is of course a case of extremely diverging beliefs. On the other hand, consider the case of belief homogeneity, $\mathbb{Q}_{1}=\dots=\mathbb{Q}_{n}=\mathbb{P}$. Then, we will have $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$ if and only if $\sum_{i=1}^{n}\lambda_{i}^{+}<1$, which reflects the impact of preferences. Clearly, when the values of $\sum_{i=1}^{n}\lambda_{i}^{+}<1$, indicating that agents accept lower total probabilities of adverse events, the inf-convolution is finite. Whereas, the higher risk tolerance of the tail events (i.e., $\sum_{i=1}^{n}\lambda_{i}^{+}>1$) enables the possibility of finding a trivially good deal for all agents, leading to an infinite inf-convolution.
@@ 3.2 Homogeneous beliefs
In this subsection, we focus on the case $\mathbb{Q}_{1}=\dots=\mathbb{Q}_{n}=\mathbb{P}$. We find more explicit expression for the inf-convolution. Let $\Lambda^{*}(x)=\sup_{y_{1}+\dots+y_{n}=x}\left(1\wedge\sum_{i=1}^{n}\Lambda_{i}(y_{i})\right)$ for all $x\in\mathbb{R}$. We say $\Lambda^{*}$ is attainable if for each $x\in\mathbb{R}$, there exists $(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$ satisfying $\sum_{i=1}^{n}y_{i}=x$ such that $1\wedge\sum_{i=1}^{n}\Lambda_{i}(y_{i})=\Lambda^{*}(x)$. For homogeneous beliefs, $\Lambda^{*}(x)$ represents the largest aggregate risk tolerance of the agents to control the tail event $\{X>x\}$ in the definition of $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$.
@@
For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, if $\Lambda^{*}$ is attainable, then we have
@@
Moreover, if additionally $x^{*}=\Lambda^{*}\mathrm{VaR}^{\mathbb{P}}(X)>-\infty$ and $\Lambda^{*}$ is right-continuous at $x^{*}$, then the optimal allocation is given by (3) with $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $1\wedge\sum_{i=1}^{n}\Lambda_{i}(y_{i}^{*})=\Lambda^{*}(x^{*})$ and $(A_{1}^{*},\dots,A_{n}^{*})$ satisfying $\mathbb{P}\left(X>x^{*},A_{i}^{*}\right)\leqslant\Lambda_{i}(y_{i}^{*})$.
@@
$A_{i}^{*}$ =$=\left\{1-\sum_{j=1}^{i}\Lambda_{j}(y_{j}^{*})<U_{X}^{\mathbb{P}}\leqslant 1-\sum_{j=1}^{i-1}\Lambda_{j}(y_{j}^{*})\right\},~{}i=1,\dots,n-1,$
$A_{n}^{*}$ =$=\left\{\left(1-\sum_{j=1}^{n}\Lambda_{j}(y_{j}^{*})\right)\wedge 0<U_{X}^{\mathbb{P}}\leqslant 1-\sum_{j=1}^{n-1}\Lambda_{j}(y_{j}^{*})\right\},$
@@
We can also remove the assumption on the attainability of $\Lambda^{*}$ in Theorem 2 to obtain an alternative expression. Let $\Lambda^{\mathbf{y}_{n-1}}(x)=\left(\Lambda_{n}(x-y_{n-1})+\sum_{i=1}^{n-1}\Lambda_{i}(y_{i}-y_{i-1})\right)\wedge 1$ for $n\geqslant 2$, where $\mathbf{y}_{n-1}=(y_{1},\dots,y_{n-1})$ and $y_{0}=0$.
@@
For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, we have
@@ 3.3 Conditional beliefs
In this subsection, we study further how the relationship among $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ can affect the inf-convolution. In particular, we consider a case where agents operate with different information. To model this scenario, consider the case $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\in\mathcal{F}$ with $\mathbb{P}(B_{i})>0$. If $\mathbb{P}$ represents the physical probability, and $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ are the beliefs of the agents, then those $\mathbb{Q}_{i}$ demonstrate that the agents’ may have different assessments of the occurrence of events due to the asymmetric information or being subject to different regulating agencies with heterogeneous stressing events; see e.g., Acciaio and Svindland (2009). For $B_{1},\dots,B_{n}\in\mathcal{F}$ and $\Lambda_{1},\dots,\Lambda_{n}\in\mathcal{H}$, if $\mathbb{P}(\cap_{i=1}^{n}B_{i})>0$, let $\Lambda^{\diamond}(x)=\sup_{y_{1}+\dots+y_{n}=x}\left(1\wedge\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\right)$ for all $x\in\mathbb{R}$. Note that $\Lambda^{\diamond}(x)$ can be viewed as the largest weighted aggregate risk tolerance of the agents to control the tail event $\{X>x\}$.
@@
Suppose $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\subset\Omega$ with $\mathbb{P}(B_{i})>0,~{}i=1,\dots,n$. For $\Lambda_{i}\in\mathcal{H}$ with $10<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<$ 1, if $\mathbb{P}(\cap_{i=1}^{n}B_{i})=0$, then
@@
where $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|\cap_{i=1}^{n}B_{i})$. Moreover, if additionally $x^{*}=\Lambda^{\diamond}\mathrm{VaR}^{\mathbb{Q}}(X)>-\infty$ and $\Lambda^{\diamond}$ is right-continuous at $x^{*}$, then the optimal allocation is given by (3) with $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}=\Lambda^{\diamond}(x^{*})$ and $(A_{1}^{*},\dots,A_{n}^{*})$ satisfying $\mathbb{P}\left(X>\sum_{i=1}^{n}y_{i}^{*},\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{i}^{*}\right)\leqslant\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})$.
Note that Theorem 3 shows that the optimal sets $(A_{1}^{*},\dots,A_{n}^{*})$ may not be unique. Under the assumption of Theorem 3, for $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}=\Lambda^{\diamond}(x^{*})$, optimal sets can be chosen as
$A_{i}^{*}$ =$=\left\{1-\sum_{j=1}^{i}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}<U_{X}^{\mathbb{Q}}\leqslant 1-\sum_{j=1}^{i-1}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\right\},~{}i=1,\dots,n-1,$
$A_{n}^{*}$ =$=\left\{\left(1-\sum_{j=1}^{n}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\right)\wedge 0<U_{X}^{\mathbb{Q}}\leqslant 1-\sum_{j=1}^{n-1}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})}\right\}$
@@
$\mathbb{Q}(A_{i}^{*})=\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})},\quad i=1,\dots,n-1,~{}\mathbb{Q}(A_{n}^{*})=1-\sum_{j=1}^{n-1}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{i=1}^{n}B_{i})},$

### v3

@@ 3.1 General beliefs
In this subsection only, we assume that $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ are general probability measures that are not required to be atomless. We first introduce the following notation. Let $\Pi_{n}(\Omega)=\{(A_{1},\dots,A_{n}):\cup_{i=i}^{n}A_{i}=\Omega,~A_{i}\cap A_{j}=\emptyset,~i\neq j\}$. For $\Lambda_{1},\dots,\Lambda_{n}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$, let
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=\inf\left\{\sum_{j=1}^{n}y_{j}:\mathbb{Q}_{i}\left(\left\{X>\sum_{j=1}^{n}y_{j}\right\}\cap A_{i}\right)\leqslant\Lambda_{i}(y_{i})~\text{for some}~(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)\right\}.$}
@@
For $\Lambda_{i}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$, we have
@@
$X_{i}=\left(X-\sum_{j=1}^{n}y_{j}^{*}\right)\mathds{1}_{A_{i}^{*}}+y_{i}^{*},~i=1,\dots,n.$ (3)
@@
Note that Theorem 1 is very general, covering Theorems 1, 2, 4 of Liu (2025) and Theorem 4 of Embrechts et al. (2020) as special cases. Compared with the results in Liu (2025), our results in Theorem 1 offer semi-explicit formulas for the inf-convolution. Later, we will specify some special relation of $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ to obtain more explicit formulas for the inf-convolution. However, it is unclear what the inf-convolution of multiple $\Lambda\mathrm{VaR}$ under general heterogeneous beliefs is, leaving it as an open problem and deserving further study in future. Moreover, Theorem 1 does not explicitly show when the inf-convolution is -$\infty$. The discussion on the finiteness of $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$ is complicated as it heavily relies on the properties of Lambda functions and the relationship between $\mathbb{Q}_{i}$; see Liu (2025) and Proposition 1 in Section 2. This will be discussed in the next two propositions. Finally, note that if all $\Lambda_{i}$ are constants, then Theorem 1 boils down to Theorem 4 of Embrechts et al. (2020), where the inf-convolution of $\mathrm{VaR}$ has been studied with belief heterogeneity.
@@
For $\Lambda_{i}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$, we have
@@
    If $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>$-$\infty$.
Note that if all $\Lambda_{i}$ are constants and $\mathbb{Q}_{1}=\mathbb{Q}_{2}=\dots=\mathbb{Q}_{n}$, then Proposition 2 offers a sufficient and necessary condition for the finiteness of the inf-convolution. In general, due to the complexity of the $\Lambda_{i}$ functions and the heterogeneity of probability measures, those conditions are sufficient but not necessary.
@@
Suppose all $\Lambda_{i}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$ are decreasing. Then we have the following conclusion.
@@
    If $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}<1$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)=$-$\infty$;
@@
    If $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$, then $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)>$-$\infty$.
Suppose all $\Lambda_{i}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$ are increasing. Then we have the following conclusion.
@@
In this more refined result, we see that $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}$ plays a pivotal role when $\Lambda$ functions are decreasing. We see that, in order to guarantee a finite inf-convolution for decreasing $\Lambda_{i}$, it is necessary for this quantity to exceed 1. For simplicity, assume that for all $i$, $\lambda_{i}^{+}$ is attained by $\Lambda_{i}$. Then, the condition in case (ii) of Proposition 3 for deceasing $\Lambda_{i}$ is violated when one can find a partition of $\Omega$ such that for all agents we have $\mathbb{Q}_{i}(A_{i})\leqslant\lambda_{i}^{+}$. This means that for each agent we can find a set $A_{i}$ on which the risk is suitably bounded by the largest acceptable probability to the agent of an adverse tail event. We can consider $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}$ as a measure of belief homogeneity, which also reflects preferences. Consider two extreme cases. First, if sets $A_{1},\dots,A_{n}$ exist for which we have $\mathbb{Q}_{i}(A_{i})=0$ for all $i$, then we will have $\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}=0.$ The ability to find such sets is of course a case of extremely diverging beliefs. On the other hand, consider the case of belief homogeneity, $\mathbb{Q}_{1}=\dots=\mathbb{Q}_{n}=\mathbb{P}$. If $\mathbb{P}$ is atomless, then we will have $\inf_{(A_{1},\dots,A_{n})\in\Pi_{n}(\Omega)}\bigvee_{i=1}^{n}\frac{\mathbb{Q}_{i}(A_{i})}{\lambda_{i}^{+}}>1$ if and only if $\sum_{i=1}^{n}\lambda_{i}^{+}<1$, which reflects the impact of preferences. Clearly, when the values of $\sum_{i=1}^{n}\lambda_{i}^{+}<1$, indicating that agents accept lower total probabilities of adverse events, the inf-convolution is finite. Whereas, the higher risk tolerance of the tail events (i.e., $\sum_{i=1}^{n}\lambda_{i}^{+}>1$) enables the possibility of finding a trivially good deal for all agents, leading to an infinite inf-convolution.
@@ 3.2 Homogeneous beliefs
In this subsection, we focus on the case $\mathbb{Q}_{1}=\dots=\mathbb{Q}_{n}=\mathbb{P}$. We find more explicit expression for the inf-convolution. For $x\in\mathbb{R}$, let
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Lambda^{*}(x)=\sup_{y_{1}+\dots+y_{n}=x}\left(1\wedge\sum_{i=1}^{n}\Lambda_{i}(y_{i})\right)$}.
We say $\Lambda^{*}$ is attainable if for each $x\in\mathbb{R}$, there exists $(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$ satisfying $\sum_{i=1}^{n}y_{i}=x$ such that $1\wedge\sum_{i=1}^{n}\Lambda_{i}(y_{i})=\Lambda^{*}(x)$. For homogeneous beliefs, $\Lambda^{*}(x)$ represents the largest aggregate risk tolerance of the agents to control the tail event $\{X>x\}$ in the definition of $\Gamma_{\Lambda_{1},\dots,\Lambda_{n}}(X)$.
@@
For $\Lambda_{i}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$, if $\Lambda^{*}$ is attainable, then we have
@@
Moreover, if additionally $x^{*}=\Lambda^{*}\mathrm{VaR}^{\mathbb{P}}(X)>$-$\infty$ and $\Lambda^{*}$ is right-continuous at $x^{*}$, then the optimal allocation is given by (3) with $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $1\wedge\sum_{i=1}^{n}\Lambda_{i}(y_{i}^{*})=\Lambda^{*}(x^{*})$ and $(A_{1}^{*},\dots,A_{n}^{*})$ satisfying $\mathbb{P}\left(\{X>x^{*}\}\cap A_{i}^{*}\right)\leqslant\Lambda_{i}(y_{i}^{*})$.
@@
$A_{i}^{*}$ $=\left\{1-\sum_{j=1}^{i}\Lambda_{j}(y_{j}^{*})<U_{X}^{\mathbb{P}}\leqslant 1-\sum_{j=1}^{i-1}\Lambda_{j}(y_{j}^{*})\right\},~i=1,\dots,n-1,$
$A_{n}^{*}$ $=\left\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(1-\sum_{j=1}^{n}\Lambda_{j}(y_{j}^{*})\right)\vee 0}<U_{X}^{\mathbb{P}}\leqslant 1-\sum_{j=1}^{n-1}\Lambda_{j}(y_{j}^{*})\right\},$
@@
We can also remove the assumption on the attainability of $\Lambda^{*}$ in Theorem 2 to obtain an alternative expression. For $n\geqslant 2$, let
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Lambda^{\mathbf{y}_{n-1}}(x)=\left(\Lambda_{n}(x-y_{n-1})+\sum_{i=1}^{n-1}\Lambda_{i}(y_{i}-y_{i-1})\right)\wedge 1$},
where $\mathbf{y}_{n-1}=(y_{1},\dots,y_{n-1})$ and $y_{0}=0$.
@@
For $\Lambda_{i}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$, we have
@@
Different from Theorem 1, the assumption of measures being atomless cannot be removed in Theorem 2. This can be seen from the following example.
Example 1.
For $m\geqslant 2$, suppose there exist distinct $w_{1},w_{2},\dots,w_{m}\in\Omega$ such that $\{w_{1}\},\dots,\{w_{n}\}\in\mathcal{F}$ and $\mathbb{P}({w_{1}})=\dots=\mathbb{P}({w_{m}})=1/m$. Define a random variable $X$ with $X(w_{i})=i,~i=1,\dots,m$ and $X(w)=0$ if $w\notin\{w_{1},\dots,w_{m}\}$. Clearly, $X$ is measurable under $(\Omega,\mathcal{F})$. Moreover, suppose $\Lambda_{1}=\Lambda_{2}=1/(2m)$. It follows from Theorem 1 that
$~~\Lambda_{1}\mathrm{VaR}^{\mathbb{P}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{P}}(X)$
$=\inf\{y\in\mathbb{R}:\mathbb{P}(\{X>y\}\cap A))\vee\mathbb{P}(\{X>y\}\cap A^{c}))\leqslant 1/(2m)~\text{for some}~A\in\mathcal{F}\}=m.$
On the other hand,
$\Lambda^{*}\mathrm{VaR}^{\mathbb{P}}(X)=\mathrm{VaR}_{1/m}^{\mathbb{P}}(X)=m-1.$
Hence, $\Lambda_{1}\mathrm{VaR}^{\mathbb{P}}\mathop{\square}\displaylimits\Lambda_{2}\mathrm{VaR}^{\mathbb{P}}(X)\neq\Lambda^{*}\mathrm{VaR}^{\mathbb{P}}(X)$. This implies that the conclusion in Theorem 2 may not hold if the atomlessness assumption is removed.
@@ 3.3 Conditional beliefs
In this subsection, we study further how the relationship among $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ can affect the inf-convolution. In particular, we consider a case where agents operate with different information. To model this scenario, consider the case $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\in\mathcal{F}$ with $\mathbb{P}(B_{i})>0$. If $\mathbb{P}$ represents the physical probability, and $\mathbb{Q}_{1},\dots,\mathbb{Q}_{n}$ are the beliefs of the agents, then those $\mathbb{Q}_{i}$ demonstrate that the agents’ may have different assessments of the occurrence of events due to the asymmetric information or being subject to different regulating agencies with heterogeneous stressing events; see e.g., Acciaio and Svindland (2009). For $B_{1},\dots,B_{n}\in\mathcal{F}$ and $\Lambda_{1},\dots,\Lambda_{n}\in\mathcal{H}$, if $\mathbb{P}(\cap_{i=1}^{n}B_{i})>0$, let
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Lambda^{\diamond}(x)=\sup_{y_{1}+\dots+y_{n}=x}\left(1\wedge\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})}\right)$}
for all $x\in\mathbb{R}$. Note that $\Lambda^{\diamond}(x)$ can be viewed as the largest weighted aggregate risk tolerance of the agents to control the tail event $\{X>x\}$.
@@
Suppose $\mathbb{Q}_{i}(\cdot)=\mathbb{P}(\cdot|B_{i})$ for some $B_{i}\subset\Omega$ with $\mathbb{P}(B_{i})>0,~i=1,\dots,n$. For $\Lambda_{i}\in\mathcal{H}$ with $0<\lambda_{i}^{-}\leqslant\lambda_{i}^{+}<1$, if $\mathbb{P}(\cap_{i=1}^{n}B_{i})=0$, then
@@
where $\mathbb{Q}(\cdot)=\mathbb{P}(\cdot|\cap_{i=1}^{n}B_{i})$. Moreover, if additionally $x^{*}=\Lambda^{\diamond}\mathrm{VaR}^{\mathbb{Q}}(X)>$-$\infty$ and $\Lambda^{\diamond}$ is right-continuous at $x^{*}$, then the optimal allocation is given by (3) with $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})}=\Lambda^{\diamond}(x^{*})$ and $(A_{1}^{*},\dots,A_{n}^{*})$ satisfying $\mathbb{P}\left(\{X>\sum_{i=1}^{n}y_{i}^{*}\}\cap\left(\cap_{i=1}^{n}B_{i}\right)\cap A_{j}^{*}\right)\leqslant\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})$.
Note that Theorem 3 shows that the optimal sets $(A_{1}^{*},\dots,A_{n}^{*})$ may not be unique. Under the assumption of Theorem 3, for $(y_{1}^{*},\dots,y_{n}^{*})$ satisfying $\sum_{i=1}^{n}\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})}{\mathbb{P}(\cap_{j=1}^{n}B_{j})}=\Lambda^{\diamond}(x^{*})$, optimal sets can be chosen as
$A_{i}^{*}$ $=\left\{1-\sum_{j=1}^{i}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{k=1}^{n}B_{k})}<U_{X}^{\mathbb{Q}}\leqslant 1-\sum_{j=1}^{i-1}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{k=1}^{n}B_{k})}\right\},~i=1,\dots,n-1,$
$A_{n}^{*}$ $=\left\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(1-\sum_{j=1}^{n}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{k=1}^{n}B_{k})}\right)\vee 0}<U_{X}^{\mathbb{Q}}\leqslant 1-\sum_{j=1}^{n-1}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{k=1}^{n}B_{k})}\right\}$
@@
$\mathbb{Q}(A_{i}^{*})=\frac{\mathbb{P}(B_{i})\Lambda_{i}(y_{i}^{*})}{\mathbb{P}(\cap_{k=1}^{n}B_{k})},\quad i=1,\dots,n-1,~\mathbb{Q}(A_{n}^{*})=1-\sum_{j=1}^{n-1}\frac{\mathbb{P}(B_{j})\Lambda_{j}(y_{j}^{*})}{\mathbb{P}(\cap_{k=1}^{n}B_{k})},$